OpenStudy (anonymous):

FAN AND MEDAL: You decide to put $150 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2.5% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds? 10.1101 years 119.954 years 121.321 years 134.34 years

OpenStudy (anonymous):

So how do we plug them in?

OpenStudy (kash_thesmartguy):

Sorry, never mind, I was wrong

OpenStudy (anonymous):

How do I do this then? :(

OpenStudy (kash_thesmartguy):

Here- this is the formula for compound interest. A = P (1 + r/n) ^ nt

OpenStudy (anonymous):

So how do I plug in the values?

OpenStudy (kash_thesmartguy):

Sorry slove for T. T is your answer. A=150(1+0.025/12)^12(t)

OpenStudy (kash_thesmartguy):

Sorry for the mistakes. I'm new to the topic!

OpenStudy (anonymous):

A=0.1325^12(t) right?

OpenStudy (kash_thesmartguy):

How did you get that?

OpenStudy (kash_thesmartguy):

@jim_thompson5910

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Kash_TheSmartGuy I think I did it wrong the answer is A right>

jimthompson5910 (jim_thompson5910):

\[\Large FV = PV*\left(1+\frac{r}{n}\right)^{n*t}\] \[\Large 3000 = 150*\left(1+\frac{0.025}{12}\right)^{12*t}\] solve for t

OpenStudy (kash_thesmartguy):

No idea! Completely new ti the topic.

OpenStudy (anonymous):

Can someone take me step bty step I dont know how to solve for t

OpenStudy (kash_thesmartguy):

Ok now I know it..

OpenStudy (kash_thesmartguy):

First solve the brackets.. division first. Divide 0.025 by 12

OpenStudy (anonymous):

480

OpenStudy (kash_thesmartguy):

What? No, it's 0.00208333333

OpenStudy (kash_thesmartguy):

Then add 1 to 0.00208333333

OpenStudy (kash_thesmartguy):

That would be 1.00208333333

OpenStudy (anonymous):

Oh my bad, I did 12/0.025 okay so what is the next step?

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

@alexistheking777 , sorry for that.

OpenStudy (kash_thesmartguy):

3000=150*1.00208333333*12^t

OpenStudy (kash_thesmartguy):

Can you do it yourself?